RXTE Helpdesk/FAQ RXTE What's New HEASARC Site Map

  • 2014 ... XTE PCA HV ON/OFF History

  • December 14 2009 - March 29 2010 HEXTE Cluster B

    HEXTE Cluster B ceased modulation (i.e., began "staring" instead of "rocking") on December 14, 2009, at approximately 16:10 UT. In January, the RXTE team performed a reboot of cluster B that successfully restored modulation, but only for a brief period before it ceased again. After careful consideration, the RXTE team has decided to permanently leave HEXTE Cluster B in an off-source position, rather than risk it becoming unrecoverable in some other position. The detectors in cluster B are still functioning well and collecting data. The off-source position of data taken by HEXTE Cluster B will thus complement the data taken by HEXTE Cluster A, which, since October 20, 2006, has been "staring" in an on-source position. Software is available to allow HEXTE Cluster A on-source data to be analyzed using HEXTE Cluster B background data.

    HEXTE Cluster B Update

  • December 25 2006 PCU 1 Loses Top Propane Layer

    On Dec. 25 at 12:56:08 UT, PCU 1 (in the 0-4 numbering scheme) developed a pinhole leak in its propane layer, consistent with a micrometeorite hit. This is similar to the loss PCU 0 suffered in May 2000 and in keeping with estimated micrometor hit rates. The impact on performance of the entire PCA is expected to be minimal; PCU 1 is only used about 15% of the time and not at all for weak sources. It's performance should be similar to that of PCU 0 now, suitable for use on galactic sources and timing observations. The PCA team cautions that a new response matrix and possibly new background model will be necessary for PCU 1; work is beginning on those issues, and expected to take several months. PCU 1 was turned off when the decreasing propane pressure was detected, and turned back on at 20:29 UT on Dec. 28. RXTE Standard Products lightcurves and spectra from Dec. 28 on will not include data from PCU 1. (Just as Standard Products made after the May 2000 PCU 0 propane loss do not include PCU 0 data.)

  • October 20 2006 HEXTE Cluster A Rocking Anomaly

    Earlier this year, HEXTE cluster A experienced more periods where on- and off-source modulation ceased. As this behavior became more frequent, the decision was made to permanently leave HEXTE Cluster A in the on-source position, rather than risk it becoming unrecoverable in an off-source position. The detectors in cluster A are still functioning well and collecting data; HEXTE cluster B also still modulates as normal, and shows no sign of impairment. The HEXTE team has delivered new software that will allow users to generate cluster A background estimates from simultaneous cluster B background files (see Katja Pottschmidt's HEAD poster about hextebackest). The RXTE GOF is now working to make this code available at the next interim HEAsoft (FTOOLS) release, currently scheduled for late November. Before analyzing any HEXTE A data taken in 2006, we recommend users consult the RXTE "Significant Events" page to see whether HEXTE A was fixed or rocking when their data was collected.

    When the HEAsoft release containing the new HEXTE tool is available, an announcement will be sent to the XTENEWS mailing list, and will also appear on the RXTE homepage. If you have any questions, please send them to the RXTE GOF using our Feedback form.

     Jan. 25 (2006:025) 09:59:59 UT - cluster A stopped rocking
     Jan. 27 (2006:027) 21:02:40 UT - rebooted, rocking resumed
     
     Jan. 31 (2006:031) 01:26:27 UT - cluster A stopped rocking
     Jan. 31 (2006:031) 15:50:19 UT - rebooted, rocking resumed
     
     Feb. 14 (2006:045) 08:15:30 UT - cluster A stopped rocking
     Feb. 14 (2006:045) 21:27:30 UT - rebooted, rocking resumed
     
     Feb. 16 (2006:047) 21:33:55 UT - cluster A stopped rocking
     Feb. 17 (2006:048) 00:23:03 UT - rebooted, rocking resumed
     
     Feb. 22 (2006:053) 22:16:02 UT - cluster A stopped rocking
     Feb. 24 (2006:055) 00:19:04 UT - rebooted, rocking resumed
     
     Feb. 25 (2006:056) 16:17:04 UT - cluster A stopped rocking
     Feb. 27 (2006:058) 19:01:05 UT - rebooted, rocking resumed
     
     Mar.  6 (2006:065) 13:21:07 UT - cluster A stopped rocking
     Mar.  6 (2006:065) 15:59:32 UT - rebooted, rocking resumed
     
     Mar.  7 (2006:066) 15:17:53 UT - cluster A stopped rocking
     Mar.  7 (2006:066) 16:14:28 UT - rebooted, rocking resumed
     
     Mar.  9 (2006:068) 17:34:35 UT - cluster A stopped rocking
     Mar.  9 (2006:068) 19:49:49 UT - rebooted, rocking resumed
     
     Mar. 14 (2006:073) 20:13:52 - stopped rocking at on-source position; 
     			    not rebooted (left in on-source position)
     
     May  22 (2006:142) 11:01:36  - system unexpectedly restarted itself,
                                    tried to rock, froze in off-source position
     
     Jul. 13 (2006:194) 16:48:48 - cluster A moved to on-source position
                        16:49:52 - brief period of rocking begins
                        16:54:08 - on-source continuous stare begins
                        16:56:48 - cluster A rebooted, resumed continuous
                                   on-source stare after that; modulation
                                   disabled from now on, to prevent any
                                   future move from on-source position
    
    

  • December 12 2005 - January 4 2006 HEXTE Cluster A Rocking Anomaly

    Between Dec. 12, 2005 01:26:24 UTC and Jan. 4, 2006 22:40:40 UTC HEXTE cluster A again did not modulate on and off source. As before, the detectors were on, but the cluster position remainedfixed for the entire period. Cluster B functioned normally throughout. This time cluster A stopped close to one of the background positions with little exposure to the source. Again a reboot of cluster A and reinitialization of the rocking has restored full function. We believe it is a coincidence that this anomaly occurred very close to one year after the previous anomaly.

  • December 13 2004 - January 14 2005 HEXTE Cluster A Rocking Anomaly

    Between Dec. 13, 2004 and 18:37:40 on Jan. 14, 2005, HEXTE cluster A did not modulate on and off source. The detectors were on, but the cluster position remained fixed for the entire period. Cluster B functioned normally throughout. Analysis of engineering data taken while cluster A was not modulating led to the conclusion that sensors responsible for detecting cluster position had become confused, triggering a shutdown of the software controlling cluster motion. A reboot of the relevant systems on January 14 restored full function for cluster A. There is no evidence of any hardware or software damage, and the cluster is behaving normally. Data taken during the "fixed position" period indicates cluster A stopped as it was rocking between the -1.5 degrees off-source position and its full on-source position. An analysis of a V0332+53 observation by Wayne Coburn found that cluster A had 75% of normal response to on-source targets during this period. The HEXTE team intends to provide boresight and other information to aid in analysis of cluster A data from this period. Further details will be posted on the RXTE homepage, as they become available.

  • July 25 2004. Incorrect Command Load

    An incorrect command load for half of July 25 2004 was sent to RXTE. This resulted in several observations not being carried out but an extended observation of 4U 1702-42 took place instead. There was also a loss of good time with several PCUs not being turned on for part of this period. The observations not carried out were:

    90097-01-51-20 315.8880  45.7500   0.70 04/07/25 14:09  SAX_J2103.5+4545 
    90147-01-50-00 311.0410 -10.7230   1.36 04/07/25 14:44  MARK_509 
    90801-05-20-04 345.0000 -18.0000   0.48 04/07/25 15:17  BACKGROUND#5 
    90056-02-22-00 272.1170 -36.9780   7.64 04/07/25 15:36  GALACTIC_BULGE_2 
    80033-01-19-04 256.5630 -43.0360   3.46 04/07/25 19:04  4U_1702-42
    60705-01-67-00 255.7060 -48.7900   0.94 04/07/25 20:04  GX_339-4 
    90410-01-15-00 248.5070 -47.3930   0.94 04/07/25 20:33  4U_1630-47 
    90107-01-03-02 274.8300 -25.4100   0.92 04/07/25 20:50  V_4641_SGR
    90152-02-19-00  68.2960   5.3540   1.50 04/07/25 21:12  3C_120
    90026-01-04-00 299.5900  35.2020   3.48 04/07/25 22:05  CYG_X-1
    

  • March 31 - April 1 2003. HEXTE Anomaly
    At 17:21:19 UT on 31 March 2003, HEXTE cluster A apparently sustained a single event upset in its telemetry interface. This caused its data stream to be corrupted, although the cluster apparently continued to operate normally. On 1 April at 18:25, the cluster was successfully rebooted. Note that observations between 18:25 and 22:39 have non-standard science mode and rocking parameters. Normal science operations resumed at 22:39 when the spacecraft schedule configured the cluster for observation 70078-01-33-00. Observations made between 2003:03:31 at 17:21 and 2003:04:01 22:39 were affected. Cluster B continued to operate nominally during this period.

  • November 25 2001. ACS Problem
    Due to an anomaly with the attitude control system slews were disabled.

  • May 17 2001. TDRSS Contact Schedule Problem
    The CES (Contact Event Schedule for TDRSS) load onboard XTE lapsed. At the beginning of May 17th there was no active schedule aboard the spacecraft. The problem was fully resolved by the TDRSS contact that started at 13:38. The lack of TDRSS contacts during this period resulted in the data recorders on board the spacecraft being overwritten with the loss of some spacecraft and housekeeping data.

  • September 6 - 13 2000. Attitude Errors
    Between September 6 16:00 to September 13 15:40 (UT) 2000, there were errors in the spacecraft attitude associated with the star trackers not locking onto stars. All data during this period may be subject to attitude errors of up to 1 degree which will not be reflected in the attitude file. PCA and HEXTE data obtained during this period can therefore be subject to very large errors on the measured flux.

  • May 12 2000 PCU 1 (0) Problems
    Starting at about 1:06 UT May 12, 2000, the pressure in the propane volume of PCU 0 suddenly began decreasing. The propane pressure is expected to reach zero within a day. The PCU 0 high voltage is currently off. At this moment, we expect the xenon volume will be maintained, and we will test this early next week after the pressures have become stable. We expect that PCU 0 will remain functional, have an increased background rate, and a somewhat different gain. We will provide additional information next week. - Keith Jahoda for the PCA team

  • March 5 2000 Spacecraft Data Loss Due to RAID Failure
    Due to a RAID failure in the MOC, space was not released on the on-board data recorder. This resulted in the loss of spacecraft data (e.g. attitude information) although science data was unaffected and should still be usable. Sources affected are GX5-1, SS Cyg, XTE J1543-56 and a background observation.

  • September 13 1999 Command Load Problem
    An incorrect command load was uplinked to the spacecraft for the period starting 13:44. The spacecraft detected the problem and therefore no commands were executed by the spacecraft until the correct command load was uplinked at 16:20. The High Rate Monitor turned off the PCA after it entered the SAA at 14:13 and it remained off until the SAA exit at 16:28. The ASM HV was turned off until the start of the next day.

  • September 8 1999 TDRSS/RXTE Communication Problem
    On September 9 several scheduled communcation contacts between TDRSS satellites and RXTE had problems. Because of this there was some loss of data. The problem is thought to be due to a reduced accuracy for the ephemeris for the orbit of RXTE. The ephemeris problems are likely indirectly associated with the transmitter problem and improved ephemerides are now being produced.

  • September 1 1999 Spacecraft Transmitter Problem
    On September 1 one of the two transmitters on board RXTE suddenly stopped working. The effect of this is that currently only one of the two high gain antennas (HGA) carried by RXTE can now be used. The problem is being investigated and additional high rate downlinks are being scheduled to compensate for this loss of an HGA and avoid data loss. There will, however, now be longer gaps during which satellite telemetry is not received.

    On Septemer 1 (day 244) some loss of data did occur:

    APID   TIME       NUMBER OF MISSING PACKETS
    
    55   18:52-19:39       194
    59   18:52-19:39       184
    63   18:52-19:39       725
    70   02:10:02:31       198
    70   03:00-05:56       814
    70   19:02-19:39       374
    74   02:12-02:31       572
    74   03:01-05:52      2377
    74   19:03-19:39      1064
    79   19:03-19:39       719
    80   01:55-02:33      2015
    80   20:12-21:48      4224
    86   01:55-02:33      1925
    86   20:12-21:48      4137
    
    There was no loss of data on day 245.

  • March 22 1999 PCU Gain Reduction
    On March 22 at around 16:30 new onboard procedures (RTSs) were loaded up to RXTE to run the PCUs at the reduced high voltage setting of "4".

  • March 23-25 1999 PCU 4/5 Gain Anomalies
    Between March 23-25 PCU 4 and 5 operated for some periods at the previous high voltage setting of 5 rather than 4. This was due to incorrect settings in on-board procedures called after a scheduled rest periods for these detectors. The time periods affected were:
    1999:082:06:02 - 07:41   4U_0115+634  (PCU 4 at level 5), 40051-05-03-00 S. Zhang
            :09:23 - 11:02   CAS_A and 4U_0115+634  (PCU 4 at level 5), 40806-01-02-00, 
    40411-01-21-00, Public
            :18:06 - 19:12   XTE_J1716-389, GRB990323  (PCU 4 at level 5), 40037-01-08-00 LEVINE,
    40413-01-02-00  Public
    
        :083:12:42 - 13:22   4U_0115+634  (PCU 4 at level 5), 40411-01-22-00 Public
            :18:22 - 19:06   XTE_J1550-564  (PCU 4 and 5 at level 5), 40401-01-64-00 Public
    
        :084:06:00 - 07:39   4U_0115+634 and Background (PCU 4 and 5 at level 5), 40051-05-04-00
    S. Zhang,  40801-06-12-05 Public
    
    

  • March 17 1999 Temporary PCU Gain Reduction
    On 1999 March 17 from 03:05 UT to 04:08 all 5 PCUs were temporarily operated at a reduced high voltage setting during a calibration observation of the Crab. This was to investigate the effects of operating at a reduced HV level on detector performance.

  • March 14 - 18 1999 Bad memory card in virtual recorder
    Between March 14, 7 UT and March 18, 18 UT RXTE was operating with a faulty memory card that led to some corrupted data. The problem was isolated to half of one of seven cards and the memory was remapped around that area (7% of the total) on March 18 and no further data corruption should occur due to this problem after that date.

  • March 1999 PCU 2
    Between the period March 9-10 PCU 2 suffered 4 breakdown events. The problem is being evaluated and PCU 2 ON/ON times are also included in the XTE PCA HV ON/OFF History

  • Oct. 19 - Nov. 5 1998 PCU 4/5 TSM
    It was discovered that the onboard automatic procedures that turn off PCUs 4 and 5 when possible breakdowns occur were not fully re-enabled after the earlier spacecraft computer crash. A result is that data from PCU 4 and 5 may be contaminated by break downs and care should be taken in analyzing data from these PCUs obtained during this period.

  • Oct. 17 -19 1998 Spacecraft computer SEU
    A single event upset (SEU) occurred in the spacecraft computer on October 17th. Full recovery occurred on October 19th. During this period there were no PCA observations and only very limited HEXTE data on X0614+091 was obtained.
  • Apr. 22, 1998. 17:30 Upload of VRTX ISR EEPROM patch
    The XTE VRTX ISREEPROM patch was successfully loaded to the spacecraft by 17:30 April 22, 1998 (UT). The spacecraft was slewed to the sunline by 15:35 and the EEPROM load was completed by 16:30. The ACS processor was then warm started to load the patch to RAM and the remaining RAM patches were reloaded by 17:30.

  • Feb.13, 1998. 20:45:35 Missed ATS buffer upload
    Due to missing ATS buffer upload, the spacecraft TSM 69 detected no buffer. ASM & PCA HVs were turned off. RXTE recovered to the normal operation as of the out-SAA at 21:43. No planned slews were missed.

  • Sep.23, 1997, 02:30 RXTE resumed to normal observation
    After the Safehold due to a cold restart of the ACS during SAA at 07:58 Sep 22. RXTE resumed to normal observation plan.

  • Jun.02, 1997, 22:48 RXTE resumed to normal observation
    After the ACS hold due to an incorrect slew of one solar array, RXTE resumed to normal observation plan.

  • May.31, 1997, 01:49 Solar Array Problem
    Due to a problem on Solar Array, all SLEWs are disabled. RXTE has been pointing on BY Cam with mainly 3 PCUs and HEXTE until the solar array problem was fixed.

  • May.15, 1997, 15:02:22...All PCUs tripped off.
    All 5 PCUs were shut off by the HRM due to large rates in the propane layer, but turned back on within a few hours. Everything appears normal.

  • Jan.15, 1997, 14:28:04...ASM recovered from Safehold
    ASM HV back on via RTS call and back to normal operations after ACE safehold.

  • Jan.15, 1997, 02:00...Spacecraft/PCA/HEXTE/DAP recovered from Safehold
    Spacecraft recovered from ACE safehold. PCA, HEXTE, EDS configured for normal operations after ACE safehold.

  • Jan.14, 1997, 19:55:45...RXTE went into ACE Safehold
    Spacecraft went into ACE Safehold mode during SAA. DAP halted and the instruments are 'safed' by safing RTS's.

  • Jan. 10-11, 1997...XTE affected by Magnetic Cloud Event (?)
    A huge magnetic cloud event in the Solar Wind hit the local Earth environment. RXTE may have effected by this event.

  • Nov. 28, 1996...EA_3 problem
    03:06:00 EA 3 remained IDLE instead of being in SB_125us_0_13_1s mode.
    03:25:46 EA 3 was commanded into SB_125us_0_13_1s mode via realtime command.

  • Sep. 23, 1996...RXTE recovered from safehold mode.
    Thanks to the efforts of spacecraft team, RXTE was successfully reconfigured to normal science mode. The DAP began execution at 267/03:55(UT). RXTE slewed to Cen X-3 at 04:16 to start observation 10134-01-02-01.

  • Sep. 22, 1996...RXTE had fallen into safehold mode and recovered.
    At 10:48am Sep. 22 (UT), due to the suspension of the communication between the onboard ACS processor and ACE hardware for several seconds during SAA passage, RXTE fell into safehold mode.

  • Sep. 18, 1996...on HEXTE from UCSD
    HEXTE deadtime counter configuration changed at UT 1996 March 28 22:10:00
    For details, see HEXTE deadtime description

  • Jul. 10, 1996...XTE Standard Roll Bias vs Sun angle
    GIF Plot / Ascii Dump / PS file

  • Jul. 5, 1996...RXTE PCA gain change history
    Startgain setting 77778
    Jan. 18, 1996Disable gain and offset because of flight software error
    Feb. 27 19:24, 1996Upload new flight software and re-enable gain/offset correction
    Mar. 25 21:00, 1996Change HV setting to 6666
    Apr. 15, 1996Change HV setting to 55555

  • July. 3, 1996...HEXTE status report
    HEXTE Cluster A experienced commanding upset

  • Apr. 12, 1996...XTE Standard Roll Bias History
    Day of 1996 Standard Roll Bias
    (degree)
    0 - 81.0 0.
    81.0 - 82.0 -4.5
    82.0 - 86.33 0.
    86.33 - 9.35 -4.5
    93.5 - 97 0.
    98 - 102.77 -8.
    102.77 - 103 -4.5
    Current Value -8.

  • [Upcoming - modification to SAA RTSs]

  • Mar. 25, 1996...PCA/EDS gain/offset uploaded

  • Went to positions closer to Sun to warm PCA
    Two of the five PCUs (4 & 5) exhibited anomalous behavior on March 19th. While the problems are being analyzed to ensure the safety of the detector the high voltage was turned off to all five PCUs. Possible solutions include adjusting spacecraft attitude slightly so as to keep the detectors warmer and turning off (rather than reducing) the high voltage during SAA passages.

  • SSC wire burn outs - SSC2 and 3
    Real Time Command History for ASM control is available. This can be used for checking the good time of each SSC.
    For more information, please consult the ASM status report which is updated with the latest developments.

  • Mar. 6, 1996...Failure of HEXTE detector.
    The HEXTE instrument team at UCSD has discovered that the pulse height analyzer for the detector #3 (out of 4) in cluster B of the HEXTE on the Rossi X-ray Timing Explorer failed on 1996 March 6 at 11:27:12 UT. The failure mode is such that the value of the pulse height in telemetry is no longer scientifically useful for this detector. However, the detector can still be used for timing studies. For more details, please see the HEXTE status report from UCSD

  • Feb. 28 15:10(ET), 1996...Name change to Rossi Explorer
    see also NASA press release (28/02/96)

  • Changed RTEs

  • Early Feb., 1996...Changes to Standard Modes
    New standard modes are now running: These enable bright sources to be observed without bin rollover.

  • Boresight correction number 1
    The "Science Axis" in spacecraft coordinates was set to (1, 0.0, 0.00070)


  • For questions about scheduling of RXTE observations please send email to xteplan@athena.gsfc.nasa.gov If you have other questions about RXTE, please send email to one of our help desks.

    This page is maintained by the RXTE SOF and was last modified on Wednesday, 01-Dec-2010 11:07:26 EST.

    Responsible NASA Official: Phil Newman
    Web Curator: Robin Corbet

    Privacy Policy and Important Notices.